There is a proton and antiproton that collide head-on with equal kinetic energies. Two gamma rays with wavelengths of 1 fm are produced. Calculate the Kinetic energy of the proton.
Answer #1

2K.E. = hc / lambda

2 x K.E = 6.63 x 10-34 x 3 x10^8 / (1 x 10-15 )

K.E. = 9.945 x 10-11 J
Answer #2

KE = h*c/lamda

= 6.625*10^-34*3*10^8/1*10^-15

= 1.9875*10^-10 J

= 1.24*10^9 eVolts